2 Please answer the following questions using the structure provided (Pr fol). Please if not applicable, write N/A ) mmber of sigma (a) bonds ANS l) number of sp hybridized atoms ANS ii) number of pi (T) bonds ANS CH3 QH CH3 v) molecular formula 3 Please provide the structure of 2-chloro-2-methyl-3-phenylheptane, Ha CH3 Propofol (p. 42, Klein (2e)) 年9e の 5
Answer #1

1) number of sigma bonds = 26 (All single bonds are sigma bonds.)

2) number of sp3 hybridised atoms = 7

3) number of pi bonds = 3 (All double bonds are pi bonds.)

4) number of sp2 hybridised atoms = 6 (All double bonded carbon atoms are sp2 hybridised atoms.)

Cl, CH3 structure of 2-chloro-2-methyl-3-phenylheptane Explanati on: The rootword is hept- so the number of carbon atoms in t